| ID |
Date |
Author |
Subject |
Text |
 |
|
62
|
Mon, 3.08.2026 09:20:00 |
A. Asai | 2026 Liquid Nitrogen Target run | [B]Data taking period:[/B]
(Thin) First run 48513, since 13/7/2026 14:33:02
(Thin) Last run 48589, until 22/7/2026 07:45:32
|
|
|
61
|
Fri, 31.07.2026 12:55:47 |
L. Ren | 2026 Thin Beryllium Target Run | Data taking period:
First run 48670, since 7/29/2026 21:33:00
Last run 48700, ended at 8/3/2026 6:02:00 |
|
|
60
|
Wed, 1.04.2026 17:09:20 |
B. Maksiak | 2026 Liquid Hydrogen Target run | [B]Data taking period:[/B]
First run 48131, since 16.03.2026 12:07 (Unix:
1773659220)
|
|
|
59
|
Mon, 1.12.2025 14:48:44 |
M. Bajda | PbPb 2025 - Open Charm Run | [B]Questions to Detector Supervisor:[/B]
[LIST]
[*] Save detector settings to elog from the |
|
|
58
|
Wed, 5.11.2025 14:54:20 |
B. Maksiak | 2025 Liquid Hydrogen Target run | [B]Target:[/B]
Liquid hydrogen target [URL=https://edms.cern.ch/document/3293065/1]Schematics[/URL],
[URL=https://edms.cern.ch/document/3365216/1]Description[/URL] |
|
|
57
|
Tue, 1.07.2025 05:04:48 |
K. Wojcik | Run settings for oxygen run 2025 | Beam: oxygen 150 A GeV/c
Target: H2O
|
|
|
56
|
Tue, 3.06.2025 09:50:30 |
K. Wojcik | Neutrino run 2025 | For detailed information look there: https://na61-elog.cern.ch/Beam+logbook/276 |
|
|
Draft
|
Mon, 2.12.2024 08:14:55 |
N. Amin | Fragmentation Run | Trigger Settings:
T1 = S1 * S2 * V1bar * S1_splitbar [charge |
|
|
Draft
|
Mon, 2.12.2024 07:54:16 |
N. Amin | Run settings for 2024 fragmentation run | Trigger settings:
T1 = S1 * S2 * V1bar * S1_splitbar (charge |
|
|
53
|
Mon, 2.12.2024 04:01:54 |
M. Bajda | 2022 Pb open charm run (ongoing - still not finished) | [B]Target:[/B]
[LIST]
[*] Pb (continuous) target (3mm Pb), space |
|
|
52
|
Mon, 2.12.2024 02:56:43 |
M. Bajda | 2023 Pb open charm run (ongoing - still not finished) | [B]Target:[/B]
[LIST]
[*] Pb (continuous) target (3mm Pb), space |
|
|
51
|
Fri, 29.11.2024 14:57:02 |
B. Maksiak, M. Bajda | 2024 Pb open charm run (ongoing - still not finished) | [B]Questions to Detector Supervisor:[/B]
[LIST]
[*] Save detector settings to elog from the |
7x |
|
50
|
Thu, 31.10.2024 16:29:45 |
B. Maksiak, M. Bajda | 2024 pi- + C | [B]Target:[/B]
[LIST]
[*] Short T2K C target 20x10 mm + 5 mm gap |
  |
|
49
|
Mon, 28.10.2024 14:58:10 |
B. Maksiak | 2024 pi+ + C | [B]Target:[/B]
[LIST]
[*] Short T2K C target 20x10 mm + 5 mm gap |
7x |
|
48
|
Mon, 8.07.2024 13:45:14 |
Y. Nagai | 2024 Neutrino run (proton 120 GeV/c on LBNF target) | Target: LBNF prototype 150 cm (Graphite type:
Toyo-Tanso IG-510, 1.9 g/cm^3 density), Al
calibration target (7cm thick, 18cm upstream |
6x |
|
48
|
Mon, 8.07.2024 13:45:14 |
Y. Nagai | 2024 Neutrino run (proton 120 GeV/c on LBNF target) | Target: LBNF prototype 150 cm (Graphite type:
Toyo-Tanso IG-510, 1.9 g/cm^3 density), Al
calibration target (7cm thick, 18cm upstream |
6x |
|
47
|
Mon, 16.10.2023 19:10:36 |
Pawel Staszel | New ALPIDE thesholds | ON 12.10.2023 evening the ALPIDE thresholds
were adjusted. It was done just after we
finished taking the geometry tuning runs |
|
|
46
|
Sun, 8.10.2023 13:51:05 |
P. Podlaski | 2023 Pb run | [COLOR=red][SIZE=6]Discovered problems by
run:[/SIZE][/COLOR]
run 43448 to 43478 no S1_1 in DRS, TOF probably |
|
|
45
|
Thu, 13.07.2023 17:04:49 |
Y. Nagai, L. Ren | Kaon + Carbon @ 60 GeV/c Run | Target: Carbon target (1+2)
Beam: kaon
|
|
|
44
|
Sun, 9.07.2023 09:28:22 |
Y. Nagai, L. Ren | Total Cross Section Run kaon + Ti @ 60 GeV/c | |
|
|
43
|
Sun, 9.07.2023 09:28:09 |
Y. Nagai, L. Ren | Total Cross Section Run proton + Ti @ 60 GeV/c | |
|
|
42
|
Sun, 9.07.2023 09:27:20 |
Y. Nagai, L. Ren | Total Cross Section Run pion + Ti @ 60 GeV/c | |
|
|
41
|
Sat, 8.07.2023 15:00:34 |
Y. Nagai, L. Ren | 60 GeV beam composition scan with full MF | Target: no target (Carbon target in OUT position)
Beam: kaon
|
|
|
40
|
Tue, 26.07.2022 11:46:25 |
Y. Nagai, M. Friend | T2K replica target run, full MF field (158 GeV field) | Target: no target (T2K replica target in OUT
position)
|
|
|
39
|
Tue, 26.07.2022 11:28:30 |
Y. Nagai, M. Friend | T2K replica target run, 1/8 MF field (20 GeV field) | Target: T2K replica 90 cm (New replica target
"C" produced in 2022)
|
|
|
38
|
Tue, 26.07.2022 10:40:55 |
Y. Nagai, M. Friend, J. Depp and A.Heard | T2K replica target run, half MF field (80 GeV field) | Target: T2K replica 90 cm (New replica target
"C" produced in 2022)
|
|
|
37
|
Fri, 30.11.2018 15:43:08 |
Lajka | Run settings for Pb+Pb @ 150A GeV/c | Target: Pb 3mm (some runs with 2mm)
[B]Beam:[/B] Pb
|
|
|
36
|
Mon, 16.07.2018 12:04:09 |
| Run settings for p+NOvA replica @ 120 GeV/c | Target: NOvA replica 130 cm (Fermilab 2018)
[B]Beam:[/B] p
|
|
|
35
|
Fri, 8.12.2017 12:18:15 |
| Run settings for Xe+La @ 13A GeV/c | Target: La 3 mm (Ki Vihao, purchased in 2014,
cut in 2017)
|
11x |
|
34
|
Fri, 8.12.2017 12:17:57 |
| Run settings for Xe+La @ 40A GeV/c | Target: La 3 mm (Ki Vihao, purchased in 2014,
cut in 2017)
|
|
|
33
|
Thu, 23.11.2017 18:18:37 |
| Run settings for Xe+La @ 30A GeV/c | On November 22 there was a risk of a power
glitch due to works on the electrical emergency
switch (AUG). We switched off magnets, and |
|
|
32
|
Mon, 13.11.2017 11:31:33 |
| Run settings for Xe+La @ 19A GeV/c |
Target: La 3 mm (Ki Vihao, purchased in 2014, |
11x |
|
31
|
Mon, 6.11.2017 20:13:52 |
| Run settings for Xe+La @ 75A GeV/c | Target: La 3 mm (Ki Vihao, purchased in 2014,
cut in 2017)
|
11x |
|
30
|
Fri, 3.11.2017 11:12:26 |
| Run settings for Xe+La @ 150 GeV/c | Target: La 3 mm (Ki Vihao, purchased in 2014,
cut in 2017)
|
11x |
|
29
|
Fri, 20.10.2017 11:53:44 |
A.Marcinek, A.Aduszkiewicz, A.Banderas | Run settings for p+Pb @ 40 GeV/c | Target: Pb 5 mm (American Elements, same as
for p+Pb in 2016)
|
11x |
|
28
|
Mon, 4.09.2017 15:52:31 |
| Run settings for neutrino physics p+C @ 90 GeV/c | Target: Carbon 14.8 mm (samples 1 and 2).
Beam: p
|
|
|
27
|
Mon, 4.09.2017 15:52:24 |
| Run settings for neutrino physics p+Be @ 120 GeV/c | Target: Beryllium 14.8 mm (samples 1 and 2).
Beam: p
|
|
|
26
|
Mon, 4.09.2017 15:52:14 |
| Run settings for neutrino physics p +C @ 120 GeV/c | Target: Carbon 14.8 mm (samples 1 and 2).
Beam: p
|
|
|
25
|
Mon, 4.09.2017 15:51:58 |
| Run settings for neutrino physics pi- +C @ 60 GeV/c | Target: Carbon 14.8 mm (samples 1 and 2).
Beam: pi-
|
|
|
24
|
Fri, 11.08.2017 12:37:13 |
| Run settings for neutrino physics pi+ +C @ 30 GeV/c | Target: Carbon 14.8 mm (samples 1 and 2).
Beam: pi+
|
11x |
|
23
|
Tue, 8.08.2017 18:40:00 |
| Run settings for neutrino physics p+Al @ 60 GeV/c | Target: Aluminium 14.8 mm (samples 1 and 2).
Beam: pi+
|
|
|
22
|
Sun, 2.07.2017 10:21:28 |
Heinrich Ignaz Franz Biber and crying Antoni | Physics data, p+Pb 30 GeV/c | Target: Pb 5 mm (American Elements same as
for p+Pb in 2016)
|
|
|
21
|
Sun, 11.12.2016 18:09:58 |
Mickey Mouse | Physics data, Pb+Pb 150A GeV/c | Target: Pb 1 mm (American Elements sample
#1)
|
|
|
20
|
Sun, 4.12.2016 11:53:12 |
S.Pulawski, E.Kaptur, A.Snoch, B. Maksiak, A.Aduszkiewicz, A.Schwarzenegger on behalf of the NA61/SHINE Collaboration | Run settings for Pb+Pb at 30A GeV/c | Target: Pb 1 mm (American Elements sample
#1)
|
|
|
19
|
Tue, 22.11.2016 17:38:28 |
S.Pulawski, E.Kaptur, A.Aduszkiewicz, Jackie Chan on behalf of the NA61/SHINE Collaboration | Run settings for Pb+Pb at 13A GeV/c | Target: Pb 1 mm (American Elements sample
#1)
|
|
|
18
|
Tue, 25.10.2016 17:34:55 |
A. Marcinek, P. Podlaski, A. Aduszkiewicz, H. Potter on behalf of the NA61/SHINE collaboration | Run settings for p+p @ 400 GeV/c | Target: LHT 20.92 mm
Beam: p
|
|
|
17
|
Sun, 9.10.2016 12:02:02 |
E. Kaptur, A. Aduszkiewicz, L. Kaczynski on behalf of the NA61/SHINE collaboration | Run settings for neutrino physics pi+Be @ 60 GeV/c | Target: Beryllium 14.8 mm (samples 1 and 2).
Beam: pi+
|
|
|
16
|
Sun, 9.10.2016 12:00:52 |
E. Kaptur, A. Aduszkiewicz, Andrey on behalf of the NA61/SHINE collaboration | Run settings for neutrino physics p+Be @ 60 GeV/c | Target: Beryllium 14.8 mm (samples 1 and 2).
Beam: p
|
|
|
15
|
Tue, 4.10.2016 10:08:12 |
E. Kaptur, A. Aduszkiewicz, H. Clinton on behalf of the NA61/SHINE collaboration | Run settings for neutrino physics p+Al @ 60 GeV/c | Target: Aluminium 14.8 mm (samples 1 and 2).
Beam: p
|
|
|
14
|
Tue, 27.09.2016 11:29:35 |
E. Kaptur, A. Aduszkiewicz, D.Trump on behalf of the NA61/SHINE collaboration | Run settings for neutrino physics p+C @ 120 GeV/c | Target: Carbon 14.8 mm (samples 1 and 2).
Beam: p
|
|
|
13
|
Wed, 21.09.2016 18:09:38 |
E. Kaptur, A. Aduszkiewicz, I.D.Know on behalf of the NA61/SHINE collaboration | Run settings for neutrino physics pi++C @ 60GeV/c | Target: Carbon 14.8 mm (samples 1 and 2).
Beam: pi+
|
|
|
12
|
Wed, 20.07.2016 17:47:54 |
M. Kuich, J. Page, E. Kaptur on behalf of the NA61/SHINE collaboration | Run settings for neutrino physics p+C @ 60GeV/c | Target: Carbon 20 mm (T2K target).
Beam: protons
|
|
|
11
|
Mon, 18.07.2016 17:21:50 |
M. Kuich, F. Mercury, E. Kaptur on behalf of the NA61/SHINE collaboration | Run settings p+Pb @ 80GeV/c | Target: Lead 5 mm
Beam: protons
|
|
|
Draft
|
Sat, 16.07.2016 17:36:22 |
A.Aduszkiewicz, M.Kuich, P. Podlaski, J.Hendrix | 2015 Basic settings for Pb @ 30 AGeV | Beamline settings:
??
|
|
|
9
|
Mon, 30.11.2015 08:27:03 |
A.Aduszkiewicz, M.Kuich, J.Morrison | 2015 Basic settings for Pb @ 30 AGeV | Beamline settings:
Beam file on beam computer
|
|
|
Draft
|
Mon, 30.11.2015 08:26:58 |
A.Aduszkiewicz, M.Kuich, J.Morrison | 2015 Basic settings for Pb @ 30 AGeV | Beamline settings:
Beam file on beam computer
|
|
|
7
|
Tue, 13.10.2015 11:50:33 |
E.Kaptur, A.Aduszkiewicz, M.Davis | 2015 Basic settings for pi+ USNA61 @ 31 AGeV | Beamline settings:
Beam file on beam computer
|
|
|
6
|
Thu, 1.10.2015 08:47:20 |
S.Pulawski | 2015 Basic settings for Ar+Sc @ 75 AGeV | Beamline settings:
Beam file on beam computer
|
|
|
5
|
Wed, 25.03.2015 15:54:42 |
A. Aduszkiewicz, S.Pulawski, B. Obama | 2015 Basic settings for Ar+Sc @ 40 AGeV | Beamline settings:
https://na61-elog.cern.ch/Beam+logbook/159
|
|
|
4
|
Tue, 17.03.2015 10:03:12 |
A. Aduszkiewicz, S.Pulawski, M.Kuich, B. Obama | 2015 Basic settings for Ar+Sc @ 30 AGeV | Beamline settings:
https://na61-elog.cern.ch/Beam+logbook/157
|
|
|
3
|
Wed, 4.03.2015 18:48:04 |
A. Aduszkiewicz, S.Pulawski, B. Obama | 2015 Basic settings for Ar+Sc @ 19 AGeV | Beamline settings:
https://na61-elog.cern.ch/Beam+logbook/156
|
|
|
2
|
Thu, 26.02.2015 13:08:44 |
A. Aduszkiewicz, S.Pulawski, B. Obama, E. Kaptur on behalf of the NA61/SHINE collaboration | 2015 Basic settings for Ar+Sc @ 13 AGeV | Beamline settings:
https://na61-elog.cern.ch/Beam+logbook/155
|
|
|
1
|
Thu, 26.02.2015 13:05:34 |
A. Aduszkiewicz, S.Pulawski, B. Obama, E. Kaptur on behalf of the NA61/SHINE collaboration | 2015 Basic settings for Ar+Sc @ 150 AGeV | Beamline settings:
https://na61-elog.cern.ch/Beam+logbook/154
|
|